A softdrink cup is cylindrical in shape. I f its volume is 1884cm3 and the radius of the circular base is 10cm, how deep is the cup?
1
Expert's answer
2012-08-10T10:37:51-0400
V = πHR² ==> H = V/(πR²) = 1884/(π*10²) ≈ 1884/(3.14*10²) = 6cm.